# FlexGuard: A Design Space for On-Body Feedback for Injury Prevention in Strength Training

### v1

@@ 2.4. Intelligent Systems in Movement and Skill Learning
In movement learning and demonstration-based systems, researchers have applied IML techniques to encode and reuse human motion, which often involves tacit knowledge that is difficult to articulate explicitly (Gillies, 2019). MotionMA (Velloso et al., 2013) enables users to demonstrate movement patterns that the system can model and compare in real time. Mimic (Mahadevan et al., 2022) applies similar principles in teleoperation, letting users record and re-trigger physical behaviors. Other approaches like danceON (Payne et al., 2021) and Rapido (Leiva et al., 2021) focus on embodied programming, through specification or demonstration, for creative and spatial applications. Vidal et al. propose a framework and infrastructure for applying IML in physiotherapy to support both in-clinic and out-of-clinic patients (Turmo Vidal et al., 2025). Haptic Scores (Escudero, 2024) attempts to map detected movements to embodied haptics for dance with EMS feedback. In sport-specific contexts, Weng et al. (Weng et al., 2025) decompose movement into feedback-friendly units to integrate AI for teaching basketball shooting skills. While recent advancements in large language models (LLMs) have significantly expanded the capabilities of intelligent systems (Mahadevan et al., 2024), designing intuitive programming environments that effectively integrate LLMs remains a challenge (Pu et al., 2025). FlexGuard builds on these directions by embedding interactive intelligence directly into the wearable interface, enabling trainees to shape feedback for safer and more personalized strength training.
@@ 3. Landscape of Feedback Modalities
Wearable systems have explored a wide range of on-body feedback modalities to support physical interaction and learning. Our review of over 50 papers from HCI research, complemented by examples from gym practices and selected work in soft robotics, identifies recurring strategies including vibration, electrical stimulation, pneumatic actuation, thermal cues, material-driven mechanisms, and structural supports. Table 1 summarizes representative work across categories.
@@
| Vibration | TrainerTap (Jo et al., 2023b), VibraForge (Huang et al., 2025), HapticHead (Kaul and Rohs, 2017), Zishi (Wang et al., 2016) | Pulses or patterns on the skin. |
| Electrical (EMS / TENS) | Teslasuit (Caserman et al., 2021), Smartwatch EMS (Takahashi et al., 2024), ReKnit-Care (Hu et al., 2025), Haptic Scores (Escudero, 2024) | Electrical stimulation of muscles or nerves. |
| Pneumatic | VabricBeads (Pardomuan et al., 2024), PneumAct (Günther et al., 2019), Force Jacket (Delazio et al., 2018), FrozenSuit (Al Maimani and Roudaut, 2017), Bellowband (Young et al., 2019), JetUnit (Zhang et al., 2024) | Air or fluid pressure for force or resistance. |
@@
| Material-Driven | Springlets (Hamdan et al., 2019), ArmDeformation (Lin et al., 2024b), ImpactVest (Tsai et al., 2022), OmniFiber (Kilic Afsar et al., 2021) | Mechanical actuation or material deformation. |
| Structural Supports | Bistable Orthoses (Lin et al., 2025b), Stiffening Garment (Kwon et al., 2022), SkinMorph (Kao et al., 2018), E-Orthosis (Yang et al., 2023), gym gear (e.g., knee sleeves (Machek et al., 2021), squat suit (Latella et al., 2024), lifting belts (Herbaut and Tuloup, 2025)) | Passive supports for posture or training. |
Across these categories, we distinguish between extrinsic mechanisms (active modalities such as vibration, electrical, pneumatic, thermal, or material-driven actuation that deliver forces or sensations to the body) and intrinsic mechanisms (passive supports such as braces, sleeves, or garments that guide or resist movement through direct material contact). Since strength training demands both corrective cues and physical support, our subsequent sections of this paper focus on these two modalities through co-design workshops exploring their affordances in training contexts.
@@ 4. On-Body Feedback for Strength Training
We focused on a set of upper-body strength exercises where errors could lead to injury, to scope the study. Although not full-body, the selected exercises encompass a broad range of movement demands in strength training, offering insights that can largely generalize to other body regions and exercises. The representative set included five exercises: dumbbell front hold, dumbbell incline fly, bench press, lateral raise, and Arnold press. Each exercise highlighted a distinct movement error risk, such as posture holding, range of motion (ROM), inter-limb balance, tempo, or complex transitions. Further details about injury risks are shown in Fig. 10.
@@ 4.1. Technology Probe Design for DIY Low-Fi Prototyping
To enable rapid DIY prototyping and immediate exploration of feedback sensations, we developed a technology probe (Hutchinson et al., 2003) designed for quick reconfiguration. The base system included adjustable shoulder braces and elbow sleeves, equipped with customizable feedback components such as inflatable air wedges on the shoulders, and rigid and stretchy straps on the elbows. The probe supported two distinct modes of on-body feedback:
@@ 4.1.1. Extrinsic Feedback
System-actuated and on-demand, simulated through the inflation of air wedges during exercise.
@@ 4.1.2. Intrinsic Feedback
Material-driven and continuously available, achieved through tension in rigid or stretchy straps that responded directly to movement.
Velcro interfaces allowed participants to quickly attach, reposition, or remove components as needed. To further support collaborative prototyping, we provided additional materials such as duct tape, resistance bands, and a mini-mannequin to facilitate communication between trainers and trainees. This setup enabled hands-on experimentation: participants configured feedback on their own bodies, performed representative movements, and reflected on the sensations in real time (Fig. 2).
@@ 4.2. Worksheet Design
We prepared a worksheet for each exercise. Each worksheet included body maps (Turmo Vidal et al., 2023) with figures of a person performing the exercise at representative moments, such as the starting position, mid-movement, and the maximum extension point. For the dumbbell front hold, an additional figure depicted an isometric hold at 15 seconds. These figures served as reference frames for participants to annotate, with the option of requesting extra sheets to sketch alternative designs or capture other parts of the movement cycle. As body maps are adaptable to diverse research contexts (Turmo Vidal et al., 2023), in our study they functioned both as a tool to guide the session and as an open-ended space for dialogue between trainer and trainee. They were designed for participants to sketch low-fidelity prototype designs, align on preferred configurations, and highlight specific body regions requiring attention. An example worksheet can be seen in Fig. 3.
@@ 4.3. Study Protocol
| Session |
Other Sports & Fitness Background |
| S1 | E1 (48, M) | 2 yrs | 5+ | – | N1 (23, M) | |||||||||
| S2 | E2 (26, M) | 1.5 yrs | 5–10 | – | N2 (26, M) | |||||||||
| S3 | E3 (25, F) | 1 yr | 2 | – | N3 (33, F) | |||||||||
| S4 | E4 (28, M) | 1 yr | 2 | – | N4 (24, M) | |||||||||
| S5 | E5 (24, F) | 8 yrs | 5+ | • Licensed Zumba & Aerobics instructor • Experience teaching classes with handheld weights | N5 (26, F) | |||||||||
| S6 | E6 (23, F) | 9 yrs | 5+ | • Division I track/cross-country athlete • Provided instruction on form and spotting to peers and juniors | N6 (23, F) | |||||||||
| S7 | E7 (24, F) | 5-7 yrs | 20+ | • National-level rowing athlete • Instructor of rowing camps | N7 (33, F) | |||||||||
| S8 | E8 (22, M) | 8 yrs | 100+ | • Fitness instructor (3 years) • Track/cross-country athlete: PAC-12 Honor Roll, National U-20 champion | N8 (25, M) | |||||||||
@@
All novice trainees (N1-N9) had less than three months of consistent strength training experience.
@@ 4.3.1. Initial Interview
Each session began with two brief, separate audio-recorded interviews, one with the trainee and one with the trainer. Participants discussed their individual training backgrounds, any prior injuries from strength training, and their experience coaching or being coached. These interviews provided context for interpreting their design decisions during the rest of the session.
@@ 4.3.2. Worksheet Introduction
DG1: Prevent injuries by providing feedback.
@@ 4.3.3. Probe Demonstration
Next, the research team introduced the technology probe. We first demonstrated the default configuration of the probe 2. Participants were then encouraged to experiment with customizing the setup to their own preferences, such as attaching straps in alternative positions (e.g., waist-to-shoulder). Both trainer and trainee wore the probe and explored three core feedback configurations:
@@ 4.3.5. Co-Design Activity and Reflection
After the initial training, each trainer–trainee pair engaged in a collaborative co-design activity by sketching their ideas on the worksheet. They then configured the probe to reflect these design decisions, after which the trainee re-performed the exercise while wearing the configured probe. Participants reflected on how the feedback felt, discussed potential improvements, and recorded notes on the worksheet. If time permitted, they iterated on the design to better align with their intentions.
@@ 4.4. Data Analysis
We conducted a thematic analysis using reflexive approaches that combined both deductive and inductive coding (Braun and Clarke, 2019) of video and audio recordings as well as annotated worksheets. The first author developed initial codes informed by theoretical frameworks from prior review articles on sports wearable technologies (e.g., (Turmo Vidal et al., 2021; Seçkin et al., 2023)), focusing on themes such as inputs, outputs, and their interplay. After collaboratively reviewing two sessions with the fourth and seventh authors, including videos (Knoblauch et al., 2014), transcripts, and worksheets, the first author generated preliminary groupings, which were then shared with the second, third, fourth, and seventh authors to refine the emerging themes. These results were iteratively discussed and revised during team meetings with the broader author group and were also compared against more recent design space publications (e.g., (Stawarz et al., 2025)). The remaining recordings were reviewed by the fourth and seventh authors under the supervision of the first author. The resulting design space was extensively discussed and refined by the first, fourth, and seventh authors, with additional input from the last author as well as the rest of the team during meetings.
5. FlexGuard: A Design Space for On-Body Feedback for Injury Prevention in Strength Training
Drawing on insights from the co-design workshops, we introduce FlexGuard, a design space that models how on-body feedback can be sensed, triggered, and delivered to support safe strength training.
@@ 5.1.1. Trigger Metric
Participants identified several physical characteristics that could serve as reliable triggers for feedback. Across multiple sessions, joint angle was the most consistently mentioned marker, particularly when arms drifted out of alignment. As E5 noted, “your elbow keeps flaring out, that’s where it starts to go wrong.” Trainers also cautioned against end-range errors—“Again, you never want to lock your elbows” (S5). Others emphasized movement dynamics (e.g., acceleration or tempo) as indicators of excessive momentum rather than controlled lifting; during a lateral raise, E4 explained to N4, “if the dumbbells bounce at the top, that’s all shoulder stress instead of control.” In S1, participants highlighted inter-limb coordination (e.g., symmetric pacing/height in bench press) as a cue to synchronize both sides. In S8, feedback was tied to muscle activation, with the trainer remarking, “make sure you engage your chest muscle, otherwise you are not working out the muscles that you want.” Three sessions (S2, S7, S9) discussed motion path, noting that a shaky bar or wobbling dumbbell trajectory signals fatigue and breakdown of posture. For instance, when E2 noticed N2 developing a slight shake during the bench press, E2 intervened to help lift the bar.
@@ 5.1.2. Trigger Policy
Participants proposed rules for deciding when feedback should be actuated and when it should be withdrawn. Two broad categories emerged. expert-based policies draw on general training knowledge, such as recommended ROM or standard joint angles. For example, most trainers emphasized enforcing limits when elbows drift too far during a press, or constraining unsafe ROM at the extremes of an incline fly. In contrast, personalized policies adapt thresholds to the individual trainee, accounting for factors such as fatigue, muscle engagement, or body type. For instance, feedback could be tuned to a user’s tolerance for fatigue or to whether the intended muscle is activating effectively, both of which can vary significantly across body types. While all policies, including expert-based, can ultimately be personalized, we make this distinction to separate those with only slight variations across most body types from those that are highly subjective.

### v2

@@ 2.4. Intelligent Systems in Movement and Skill Learning
In movement learning and demonstration-based systems, researchers have applied IML techniques to encode and reuse human motion, which often involves tacit knowledge that is difficult to articulate explicitly (Gillies, 2019). MotionMA (Velloso et al., 2013) enables users to demonstrate movement patterns that the system can model and compare in real time. Mimic (Mahadevan et al., 2022) applies similar principles in teleoperation, letting users record and re-trigger physical behaviors. Other approaches like danceON (Payne et al., 2021) and Rapido (Leiva et al., 2021) focus on embodied programming, through specification or demonstration, for creative and spatial applications. Vidal et al. propose a framework and infrastructure for applying IML in physiotherapy to support both in-clinic and out-of-clinic patients (Turmo Vidal et al., 2025). Haptic Scores (Escudero, 2024) attempts to map detected movements to embodied haptics for dance with EMS feedback. In sport-specific contexts, Weng et al. decompose movement into feedback-friendly units to integrate AI-driven textual coaching for basketball shooting (Weng et al., 2025). However, such textual feedback approaches (Mahadevan et al., 2024; Pu et al., 2025) may be less effective in physical tasks like strength training. FlexGuard builds on these directions by embedding interactive intelligence directly into the wearable interface, enabling trainees to shape feedback for safer and more personalized strength training.
@@ 3. LANDSCAPE OF FEEDBACK MODALITIES
Wearable systems have explored a wide spectrum of on-body feedback modalities to support physical interaction and learning. Our review of over 50 works from wearable computing, ubiquitous computing, and haptics research, and relevant examples from gym training practices and soft robotics surfaces several recurring strategies: vibration, electrical stimulation, pneumatic actuation, thermal cues, mechanical-based feedback, and structural supports. Table 1 summarizes representative systems and approaches across these categories.
@@
| Vibration | TrainerTap (Jo et al., 2023b), VibraForge (Huang et al., 2025), HapticHead (Kaul and Rohs, 2017), Zishi (Wang et al., 2016), HapticPilot (Sung et al., 2024) | Pulses or patterns on the skin. |
| Electrical (EMS / TENS) | Teslasuit (Caserman et al., 2021), FootStriker (Hassan et al., 2017), Smartwatch EMS (Takahashi et al., 2024), ReKnit-Care (Hu et al., 2025), Haptic Scores (Escudero, 2024) | Electrical stimulation of muscles or nerves. |
| Pneumatic | VabricBeads (Pardomuan et al., 2024), PneumAct (Günther et al., 2019), Force Jacket (Delazio et al., 2018), FrozenSuit (Al Maimani and Roudaut, 2017), aSpire (Choi et al., 2022), JetUnit (Zhang et al., 2024), OmniFiber (Kilic Afsar et al., 2021) | Air or fluid pressure for force or resistance. |
@@
| Mechanical | Springlets (Hamdan et al., 2019), ArmDeformation (Lin et al., 2024b), ImpactVest (Tsai et al., 2022) | Mechanically generated forces or deformations. |
| Structural Supports | Bistable Orthoses (Lin et al., 2025b), Stiffening Garment (Kwon et al., 2022), SkinMorph (Kao et al., 2018), ThermoFit (Wang et al., 2023), gym gear (e.g., knee sleeves (Machek et al., 2021), squat suit (Latella et al., 2024), lifting belts (Herbaut and Tuloup, 2025)) | Passive supports for posture or training. |
Across these categories, we distinguish between extrinsic mechanisms (externally actuated by components outside the wearable material, such as vibration, electrical, pneumatic, thermal, or material-driven actuation that deliver forces or sensations to the body) and intrinsic mechanisms (intrinsically produced by the wearable’s material itself; passive supports such as braces, sleeves, or garments that guide or resist movement through direct material contact). Since strength training demands both corrective cues and physical support, our subsequent sections of this paper focus on these two modalities through co-design workshops exploring their affordances in training contexts.
@@ 4. ON-BODY FEEDBACK FOR STRENGTH TRAINING
We focused on a set of upper-body strength exercises where errors could compromise safety, to scope the study. Although not full-body, the selected exercises encompass a broad range of movement demands in strength training, offering insights that can largely generalize to other body regions and exercises. The representative set included five exercises: dumbbell front hold, dumbbell incline fly, bench press, lateral raise, and Arnold press. Each exercise highlighted a distinct movement error risk, such as posture holding, range of motion (ROM), inter-limb balance, tempo, or complex transitions. Further details about safety-relevant factors are shown in Figure 10.
@@ 4.1. Technology Probe Design for DIY Low-Fi Prototyping
To enable rapid DIY prototyping and immediate exploration of feedback sensations, we developed a technology probe (Hutchinson et al., 2003; Ley-Flores et al., 2024) designed for quick reconfiguration. The base system included adjustable shoulder braces and elbow sleeves, equipped with customizable feedback components such as inflatable air wedges on the shoulders, and rigid and stretchy straps on the elbows. The probe supported two distinct modes of on-body feedback:
@@ 4.1.1. Extrinsic Feedback
Externally actuated feedback outside of the wearable’s material structure, simulated through the modulation of air wedges.
@@ 4.1.2. Intrinsic Feedback
Intrinsically generated feedback arising from the wearable’s own materials, achieved through tension in rigid or elastic straps that respond directly to the user’s movement.
We selected these modalities for the study based on their practicality for rapid prototyping in a gym environment (e.g., without electronic wiring) and their coverage of diverse experiences. The air wedges, serving as extrinsic feedback, allowed participants to explore a range of sensations on demand by varying pumping patterns and intensity. In contrast, intrinsic feedback depended almost entirely on material properties, so we provided both rigid and elastic options to offer participants a broader range of experiences.
Velcro interfaces allowed participants to quickly attach, reposition, or remove components as needed. To further support collaborative prototyping, we provided additional materials such as duct tape, resistance bands, and a mini-mannequin to facilitate communication between trainers and trainees. This setup enabled hands-on experimentation: participants configured feedback on their own bodies, performed representative movements, and reflected on the sensations in real time (Figure 2).
@@ 4.2. Worksheet Design
We prepared a worksheet for each exercise. Each worksheet included body maps (Turmo Vidal et al., 2023) with figures of a person performing the exercise at representative moments, such as the starting position, mid-movement, and the maximum extension point. For the dumbbell front hold, an additional figure depicted an isometric hold at 15 seconds. These figures served as reference frames for participants to annotate, with the option of requesting extra sheets to sketch alternative designs or capture other parts of the movement cycle. As body maps are adaptable to diverse research contexts (Turmo Vidal et al., 2023), in our study they functioned both as a tool to guide the session and as an open-ended space for dialogue between trainer and trainee. They were designed for participants to sketch low-fidelity prototype designs, align on preferred configurations, and highlight specific body regions requiring attention. An example worksheet can be seen in Figure 3.
@@ 4.3. Study Protocol
| Session | Trainer (Age, Gender) | Strength Training Experience | People Coached | Other Sports & Fitness Background | Trainee (Age, Gender) |
| S1 | E1 (48, M) | 2 yrs | 5+ | – | N1 (23, M) |
| S2 | E2 (26, M) | 1.5 yrs | 5–10 | – | N2 (26, M) |
| S3 | E3 (25, F) | 1 yr | 2 | – | N3 (33, F) |
| S4 | E4 (28, M) | 1 yr | 2 | – | N4 (24, M) |
| S5 | E5 (24, F) | 8 yrs | 5+ | • Licensed Zumba & Aerobics instructor • Experience teaching classes with handheld weights | N5 (26, F) |
| S6 | E6 (23, F) | 9 yrs | 5+ | • Division I track/cross-country athlete • Provided instruction on form and spotting to peers and juniors | N6 (23, F) |
| S7 | E7 (24, F) | 5-7 yrs | 20+ | • National-level rowing athlete • Instructor of rowing camps | N7 (33, F) |
| S8 | E8 (22, M) | 8 yrs | 100+ | • Fitness instructor (3 years) • Track/cross-country athlete: PAC-12 Honor Roll, National U-20 champion | N8 (25, M) |
@@
All novice trainees (N1–N9) had less than three months of consistent strength training experience, while expert trainers (E1–E9) were recruited based on having at least one year of consistent training experience and prior coaching experience to ensure that trainees received appropriate guidance during the study.
@@ 4.3.1. Initial Interview
Each session began with two brief, separate audio-recorded interviews, one with the trainee and one with the trainer. Participants discussed their individual training backgrounds, any prior injuries, experiences, or safety concerns related to strength training, and their experience coaching or being coached. These interviews provided context for interpreting their design decisions during the rest of the session.
@@ 4.3.2. Worksheet Introduction
DG1: Promote safety or prevent injuries by providing feedback.
@@ 4.3.3. Probe Demonstration
Next, the research team introduced the technology probe. We first demonstrated the default configuration of the probe (Figure 2A). Participants were then encouraged to experiment with customizing the setup to their own preferences, such as attaching straps in alternative positions (e.g., waist-to-shoulder). Both trainer and trainee wore the probe and explored three core feedback configurations:
@@ 4.3.5. Co-Design Activity and Reflection
After the initial training, each trainer–trainee pair engaged in a collaborative co-design activity by sketching their ideas on the worksheet. They then configured the probe to reflect these design decisions, after which the trainee re-performed the exercise while wearing the configured probe. Participants reflected on how the feedback felt, discussed potential improvements, and recorded notes on the worksheet. If time permitted, they iterated on the design to better align with their intentions. The overall DIY process is demonstrated in Figure 4
@@ 4.4. Data Analysis
We conducted a thematic analysis using reflexive approaches that combined both deductive and inductive coding (Braun and Clarke, 2019) of video and audio recordings as well as annotated worksheets. The first author developed initial codes informed by theoretical frameworks from prior review articles on sports wearable technologies (e.g., (Turmo Vidal et al., 2021; Seçkin et al., 2023)), focusing on themes such as inputs, outputs, and their interplay. Examples of early codes included “joint deviation cue” (input), “vibrotactile” (output), and “synchronous execution” (interplay), etc. After collaboratively reviewing two sessions with the fourth and seventh authors, including videos (Knoblauch et al., 2014), transcripts, and worksheets, the first author generated preliminary groupings, which were then shared with the second, third, fourth, and seventh authors to refine the emerging themes. These results were iteratively discussed and revised during team meetings with the broader author group and were also compared against more recent design space publications (e.g., (Stawarz et al., 2025)). The remaining recordings were reviewed by the fourth and seventh authors under the supervision of the first author. The resulting design space was extensively discussed and refined by the first, fourth, and seventh authors, with additional input from the last author as well as the rest of the team during meetings.
During this process, coding disagreements occasionally surfaced and were resolved through discussion. For example, when identifying the dimensions related to motion variation, one author initially preferred coding “velocity” as its own metric based on observations from lateral raise deviations, while others argued for a broader category such as “locomotion,” which could generalize to full-body exercises. Through discussion, we recognized that several correction cues extended beyond pure velocity. We therefore agreed on the more inclusive term ”movement dynamics”, which better captured the relevant changes in movement and was more appropriate for the context of this study. Another example emerged when we initially coded several dimensions using labels such as “limiting ROM extremas,” “warning,” “restricting motion path,” and “reminder.” During a team meeting, a team member noted that these labels conflated the mechanics of feedback with the user’s intended outcome. To make the dimensions agnostic to user intent, we refined them into the more neutral categories of ”opposing”, ”guiding”, ”attentional”, and ”compressive”.
5. FLEXGUARD
Drawing on insights from the co-design workshops, we introduce FlexGuard, a design space that models how on-body feedback can be sensed, triggered, and delivered to support safety scaffolding during strength training.
@@ 5.1.1. Trigger Metric
Participants identified several physical characteristics that could serve as reliable triggers for feedback. Across multiple sessions, joint angle was the most consistently mentioned marker (n=8), particularly when arms drifted out of alignment. As E3 noted, “the only thing that has to be fixed is just the angle of motion.” Trainers also cautioned against end-range errors—“People often arch their back and then tend to raise their arms. I think that might be helpful but make sure that you don’t go above range.” (S7). Others emphasized movement dynamics (e.g., acceleration or tempo) as indicators of excessive momentum rather than controlled lifting; during a lateral raise, E9 explained to N9, “Sometimes, I’m going too heavy and using too much momentum. For the last few reps, you can often try to get that extra set.” In S3, participants highlighted inter-limb coordination (e.g., symmetric pacing/height in bench press) as a cue to synchronize both sides, noted with ”It’s a problem if the right hand is further than the left hand”. In S8, feedback was tied to muscle activation, with the trainer remarking, ”…almost feel like a little bit of pressure in the spot that you’re trying to target… so you know that you’re not using the wrong muscle”. Three sessions (S2, S7, S9) discussed motion path, noting that a shaky bar or wobbling dumbbell trajectory signals fatigue and breakdown of posture. For instance, participant noted ”For this motion… If you do more up, it will hurt your joint.”
@@ 5.1.2. Trigger Policy
Participants proposed rules for deciding when feedback should be actuated and when it should be withdrawn. Two broad categories emerged. expert-based policies draw on general training knowledge, such as recommended ROM, standard joint angles, and common coaching cues. Across 6 sessions, trainers tend to transfer their expertise verbally and give interventions when trainees drifted out of widely accepted guidance. For example, most trainers emphasized enforcing limits when elbows begin to lock during bench press, ”You want to fully extend but you don’t want to lock your elbows”, or when it deviates from the shoulder-level path, ”Sometimes people will lift them like this… we don’t want that.” In contrast, across 4 sessions, participants also highlight the need for personalized policies that adapts the thresholds to individual trainee, accounting for factors such as fatigue, muscle engagement, or body type. For instance, feedback could be tuned to a user’s tolerance for fatigue or to whether the intended muscle is activating effectively, both of which can vary significantly across body types, as emphasized by E5 regarding N5’s arms length, ”Depending on how long this area is, this bubble needs to be further up as opposed to on my shoulder blade.”. While all policies, including expert-based, can ultimately be personalized, we make this distinction to separate those with only slight variations across most body types from those that are highly subjective.